# EMMA-500: Enhancing Massively Multilingual Adaptation of Large Language Models

### v1

@@ 4.11 Code Generation
Table 18: Results on Multipl-E. For language-level breakdowns, refer to Tables 26, 26 and 26 in the appendix.
@@ 6 Conclusion
This paper addresses critical advancements and challenges in adapting language models to more than 500 languages, focusing on enhancing their performance across diverse languages. We compile the MaLA corpus, a multilingual dataset for continual pre-training of multilingual language models. By expanding and augmenting existing corpora, we train the EMMA-500 model. It demonstrates notable improvements in a range of tasks such as next-token prediction, commonsense reasoning, machine translation, text classification, and open-ended generation, showing remarkable improvements in low-resource languages. Our results show that a well-curated, massively multilingual corpora can advance model capabilities. This work sets a new benchmark for inclusive and effective multilingual language models and paves the way for future research to address the disparities between high-resource and low-resource languages.
Limitations and Outlooks
@@
Our EMMA-500 model demonstrates enhanced multilingual performance relative to its base model, Llama 2 7B, as well as other continually pre-trained variants. However, it does not surpass some of the latest models, such as Llama 3, Gemma 2, and Qwen 2, on certain benchmarks. These models are likely to be a better starting point for continual pre-training compared to Llama 2 due to their larger pre-training data. We intend to pursue multilingual extension based on these newer models in the future. On the other hand, it would be practical to assess the released corpus directly in pre-training from scratch. Moreover, multilingual instruction tuning to prepare these models for better task performance and natural interactions would be useful.
@@ Acknowledgment
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546], and funding from UTTER’s Financial Support for Third Parties under the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631). The authors wish to acknowledge CSC – IT Center for Science, Finland, the Leonardo and LUMI supercomputers, owned by the EuroHPC Joint Undertaking, for providing computational resources.
Appendix A Data Sources
A.1 Monolingual Data
@@
Glot500-c uses the following datasets: AI4Bharat, [^26] https://ai4bharat.org/ AIFORTHAI-LotusCorpus, [^27] https://github.com/korakot/corpus/releases/download/v1.0/AIFORTHAI-LotusCorpus.zip Add (36), AfriBERTa (95), AfroMAFT (4; 133), Anuvaad, [^28] https://github.com/project-anuvaad/anuvaad-parallel-corpus AraBench (108), AUTSHUMATO, [^29] https://autshumato.sourceforge.net/ Bloom (66), CC100 (26), CCNet (129), CMU_Haitian_Creole, [^30] http://www.speech.cs.cmu.edu/haitian/text/ CORP.NCHLT, [^31] https://repo.sadilar.org/handle/20.500.12185/7 Clarin, [^32] https://www.clarin.si/ DART (8), Earthlings (33), FFR, [^33] https://github.com/bonaventuredossou/ffr-v1/tree/master/FFR-Dataset Flores200 (29), GiossaMedia (43; 42), Glosses (19), Habibi (35), HinDialect (13), HornMT, [^34] https://github.com/asmelashteka/HornMT IITB (61), IndicNLP (92), Indiccorp (52), isiZulu, [^35] https://zenodo.org/record/5035171 JParaCrawl (85), KinyaSMT, [^36] https://github.com/pniyongabo/kinyarwandaSMT LeipzigData (41), Lindat, [^37] https://lindat.cz/faq-repository Lingala_Song_Lyrics, [^38] https://github.com/espoirMur/songs_lyrics_webscrap Lyrics, [^39] https://lyricstranslate.com/ MC4 (103), MTData (44), MaCoCu (17), Makerere MT Corpus, [^40] https://zenodo.org/record/5089560 Masakhane community, [^41] https://github.com/masakhane-io/masakhane-community Mburisano_Covid, [^42] https://repo.sadilar.org/handle/20.500.12185/536 Menyo20K (5), Minangkabau corpora (59), MoT (97), NLLB_seed (29), Nart/abkhaz, [^43] https://huggingface.co/datasets/Nart/abkhaz_text OPUS (123), OSCAR (117), ParaCrawl (16), Parallel Corpora for Ethiopian Languages (1), Phontron (93), QADI (2), Quechua-IIC (136), SLI_GalWeb.1.0 (7), Shami (3), Stanford NLP, [^44] https://nlp.stanford.edu/ StatMT, [^45] https://statmt.org/ TICO (10), TIL (83), Tatoeba, [^46] https://tatoeba.org/en/ TeDDi (84), Tilde (107), W2C (80), WAT (91), WikiMatrix (110), Wikipedia, [^47] https://huggingface.co/datasets/wikipedia Workshop on NER for South and South East Asian Languages (114), XLSum (46). We filter out Flores200 when processing the Glot500-c dataset. Glot500-c includes texts in languages, i.e., Azerbaijani, Gujarati, Igbo, Oromo, Rundi, Tigrinya and Yoruba, from the XLSum dataset.
A.2 Code
@@
Appendix B Additional Statistics of MaLA Corpus
B.1 Supported Languages
@@ B.2 Data Analysis
Appendix C Evaluation Benchmarks
C.1 PolyWrite
The PolyWrite dataset has 51 writing tasks with the number of prompts per task shown in Figure 5. We use Google Translate to translate the English prompts into 240 languages and back-translate for translation quality assessment.
Figure 4: Writing tasks in the PolyWrite dataset.
Figure 5: Mean BLEU scores per language in the PolyWrite dataset.
Appendix D Detailed Results
@@
D.1 Commonsense Reasoning
Table 21: 0-shot results (ACC) on XCOPA in all languages

### v2

@@ 4.10 Code Generation
Table 15: Results on Multipl-E. For language-level breakdowns, refer to Tables 23, 23 and 23 in the appendix.
@@ 6 Conclusion
This paper addresses critical advancements and challenges in adapting language models to more than 500 languages, focusing on enhancing their performance across diverse languages. We compile the MaLA corpus, a multilingual dataset for continual pre-training of multilingual language models. By expanding and augmenting existing corpora, we train the EMMA-500 model. It demonstrates notable improvements in a range of tasks such as next-token prediction, commonsense reasoning, machine translation, and text classification, showing remarkable improvements in low-resource languages. Our results show that a well-curated, massively multilingual corpora can advance model capabilities. This work sets a new benchmark for inclusive and effective multilingual language models and paves the way for future research to address the disparities between high-resource and low-resource languages.
Limitations
Multilingual Benchmark
@@
Human Evaluation
While human evaluation is valuable, it comes with challenges such as subjectivity, inconsistency, and high costs, especially in a massively multilingual setting. Recruiting linguistically skilled annotators is often impractical, particularly for low-resource languages, making large-scale human evaluation unsustainable. Even assessing a subset of languages remains prohibitively expensive. While our work is constrained by these limitations, we recognize the importance of human evaluation in complementing automated metrics. However, given these challenges, we choose automatic tools rather than human evaluation to provide a more scalable and consistent alternative, despite their imperfections.
Model Performance
Our EMMA-500 model demonstrates enhanced multilingual performance relative to its base model, Llama 2 7B, as well as other continually pre-trained variants. However, it does not surpass some of the latest models, such as Llama 3, Gemma 2, and Qwen 2, on certain benchmarks. These models are likely to be a better starting point for continual pre-training compared to Llama 2 due to their larger pre-training data. We intend to pursue multilingual extension based on these newer models in the future. On the other hand, it would be practical to assess the released corpus directly during pre-training from scratch. Moreover, multilingual instruction tuning to prepare these models for better task performance and natural interactions would be useful. Despite its strong multilingual capabilities, our model still faces challenges in mathematical reasoning and machine reading comprehension. As evaluated on a machine-translated benchmark, its performance on math tasks remains limited, likely due to the inherent difficulty of numerical reasoning and the potential impact of translation artifacts on problem clarity. Similarly, while our model shows improvements in many NLP tasks, it struggles with machine comprehension, which requires deep contextual understanding and reasoning. These limitations highlight areas for future improvement, such as incorporating specialized training data or exploring alternative architectures better suited for these tasks.
Real-world Usage
This research focuses on expanding multilingual corpora and enhancing a model’s multilingual capabilities through continual pre-training. However, it is not yet ready for real-world deployment. The model has not undergone human preference alignment or red-teaming for safety and robustness. While our work makes progress in improving multilingual NLP, further refinement, including alignment with human preferences and thorough adversarial testing, would be necessary before practical deployment.
@@ Acknowledgment
The work has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101070350 and from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe funding guarantee [grant number 10052546], and funding from UTTER’s Financial Support for Third Parties under the EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631). The authors wish to acknowledge CSC - IT Center for Science, Finland, the Leonardo and LUMI supercomputers, owned by the EuroHPC Joint Undertaking, for providing computational resources.
Appendix A Questions and Answers
A.1 Novelty
Q: What is the main goal of this paper?
A: The main goal is to improve multilingual language models by creating a massive, diverse multilingual corpus (MaLA) and applying continual pre-training to enhance linguistic inclusivity and task performance.
Q: What are the main achievements and the scientific novelty?
A: Our main achievements and scientific novelty are written in the introduction (Section 1). To summarize the key contributions of this paper: the MaLA corpus’s scale, diversity, and focus on low-resource languages represent a significant advancement over previous efforts. By integrating a variety of text types—such as code, scientific papers, and instructions—this work provides a more comprehensive foundation for multilingual pre-training. These improvements lead to a better multilingual LLM (EMMA-500) that boosts multilingual capabilities evaluated on 9 tasks and 15 benchmarks. The scale and impact of this work make it a valuable contribution to the field of multilingual NLP, especially for low-resource languages.
A.2 Data
Q: How was the data mix tuned?
A: The tuning of the data mix is a separate research topic. Our paper focuses on providing a valuable resource through data processing and model training. While we did not conduct specific ablations on the data mix, we carefully curated the corpus to ensure diversity and balance across languages and text types. Further exploration of this topic could be beneficial for future work.
Q: Since whitespace cleaning only applies to whitespace-tokenized languages, are the other languages cleaner per se or need cleaning in other ways?
A: The whitespace cleaning process primarily applies to whitespace-tokenized languages, helping standardize the text for better processing. For languages that do not rely on whitespace tokenization (such as Chinese, Japanese, or Korean), the text may not require this specific cleaning but might need other preprocessing steps, such as character normalization or segmentation.
Q: How were existing datasets and their languages to be included chosen?
A: The selection of existing datasets and their languages for inclusion was based on several factors:
  • Language Coverage: We prioritized datasets that cover a wide range of languages, with a focus on including both high-resource and low-resource languages to ensure broad multilingual representation.
  • Data Quality: Datasets with high-quality, diverse content, such as books, scientific papers, and code, were selected to ensure the corpus has varied text types, improving model generalization.
  • Task Relevance: We considered datasets that aligned well with our goals for continual pre-training and multilingual adaptation across tasks like translation, classification, and reasoning.
  • Availability: Datasets that were publicly available and accessible were preferred, ensuring transparency and reproducibility.
A.3 Evaluation
Q: How do we choose the number of shots in in-context learning?
A: The choice of the number of shots in evaluation is based on the results of the Llama 2 model, which achieves reasonably good scores on high-resource languages with this setup. We do not grid-search the number of shots. We release the model weights. The readers can choose the number of shots as wish to evaluate.
Q: How would the model perform with fewer or more shots? Is the number of shots optimized for EMMA-500?
A: The choice of 3-shot in machine translation was based on results from the LLaMA 2 model, where it showed reasonable performance on high-resource languages. The number of shots was not optimized through a grid search for EMMA-500. While fewer or more shots could impact performance, further experimentation would be needed to determine the optimal number for different tasks and languages. However, this is not the primary focus of our work, which is intended as a resource paper. The 3-shot choice already effectively showcases the model’s performance across tasks, providing a good balance between data efficiency and task completion.
Q: Why no human evaluation?
A: Human evaluation, while valuable, has limitations such as subjectivity, inconsistency, and scalability. Evaluators may have biases, leading to variability in judgments across individuals and cultures. Additionally, evaluating multilingual and open-ended tasks requires diverse, linguistically skilled annotators, which can be expensive and time-consuming. These challenges make automated metrics a practical complement, even though they may lack nuanced judgment. While resource constraints have limited our ability to perform human evaluation for this work, we recognize its importance in complementing automatic metrics. However, we must emphasize the limitations and unsustainability of human evaluation, particularly in a massively multilingual setting. Even evaluating a subset of languages or datasets, most likely low-resource languages, remains prohibitively expensive. Also, finding qualified evaluators for low-resource languages is often impractical and costly. Automatic tools, while not perfect, offer a scalable and consistent method for data collection and evaluation. Determining whether automatic or human evaluation is better is not the primary focus of this paper.
Q: How reliable is machine-generated benchmark?
A: Machine-generated benchmarks have significant value as they offer scalability, consistency, and objectivity, making it feasible to evaluate models across numerous tasks and languages efficiently. They are particularly useful for comparative analysis and identifying measurable improvements. There are also limitations: they often miss nuanced qualities like creativity, contextual appropriateness, and cultural sensitivity, especially in open-ended or multilingual tasks. Biases in the generation process or evaluation metrics can also lead to misleading results.
Q: Why no safety-related evaluation?
A: We find safety-related evaluation a separate research topic. This work looks at CPT and releases a corpus and a multilingual base model rather than an instruction-tuned model. While our work focuses on multilingual data expansion and model evaluation across a broad range of languages and tasks, safety evaluation was not a primary focus as the model is not directly developed for end users. We recognize the importance of safety benchmarks like XSafety (130) and Aya redteaming (1) that only cover a few languages, and they might not fit the evaluation in a massively multilingual setting best.
A.4 Comparison with Baselines
Q: Is it unfair to compare models that do not officially support certain languages?
A: Our evaluation focuses on the performance of models across a broad range of languages, including those that may not be officially supported. This approach highlights the model’s ability to generalize and handle underrepresented languages, even if they are not specifically optimized for them. There is no widely accepted standard for which models should be used as the “fair” baseline for such comparisons.
Q: What is the relation to the MaLA-500 model/work? What are the key improvements over it?
A: The key improvements include:
  • Larger, More Diverse Corpus: our corpus includes 939 languages and 74 billion tokens.
  • Higher Token Count per Document: our corpus features longer documents (average of 90 tokens per sequence) compared to Glot500-c used by the MaLA-500 model (average of 19 tokens), providing better context and enabling the model to capture long-range dependencies.
  • More Diverse Data: We also include a wide range of text types, including code, books, and scientific papers.
  • Better Performance: Evaluated across more benchmarks, EMMA-500 shows improvements in multilingual tasks like intrinsic tasks, translation, commonsense reasoning, and text classification.
Q: Why is Tower instruct performing so poorly (especially since it is built for translation)?
A: According to its technical report, Tower is heavily optimized (through CPT and SFT) for only around 10 languages centered on translation-related data. Its language capabilities might not generalize well across low-resource languages or specific language pairs. This is essentially our model’s advantage over Tower.
A.5 Others
Q: Will this corpus, model weight, benchmark and processing scripts be public?
A: Yes, we release the MaLA corpus (including raw, cleaned, and deduplicated versions and splits for training and validation), model weights, codes, and model generation. Benchmarks, processing scripts, and training software are open-sourced by the original authors. Some benchmarks, such as PBC, will require permission to access data from the original authors.
Q: Why is the global batch size set to 16M, rather than the commonly used 4M tokens like LLaMA-2?
A: This is a technical choice that allows us to efficiently use our computing. We note that increasing batch size during training is now seen in LLM pre-training, e.g., Llama-3 doubled their batch size twice from 4M to 16M during training.
Q: How do we categorize language into different resource groups?
A: We classify languages into five groups—high, medium-high, medium, medium-low, and low—based on their token counts in the MaLA corpus. Table 17 shows the specific languages in different categories. Since the MaLA corpus is compiled from multiple corpora, its token distribution reflects the availability of textual data across languages. Grouping languages based on token counts provides a practical and data-driven way to distinguish between high- and low-resource languages, aligning with real-world data availability rather than relying on predefined taxonomies. Besides, We provide standard ISO language codes along with their corresponding writing systems, which ensures clarity, consistency, and interoperability across different linguistic resources and NLP applications.
Q: Why don’t we use Joshi et al.’s language taxonomy [^23] https://microsoft.github.io/linguisticdiversity/assets/lang2tax.txt?
A: While we acknowledge the importance of language taxonomies like Joshi et al.’s, we choose not to rely on it for this work for several reasons:
  • Focus on Practical Impact: Our work emphasizes advancing multilingual NLP through comprehensive corpus creation and improved model performance, rather than adhering to a specific taxonomy. We focused on covering as many languages as possible, particularly underrepresented ones, using ISO codes for consistency in data processing.
  • Feasibility vs. Contribution: Manually aligning our language set with Joshi’s taxonomy, especially given the absence of ISO codes, is a labor-intensive task with limited relevance to the core contributions of this paper. While it is feasible, we believe our efforts are better directed toward demonstrating the utility of the MaLA corpus and models through tangible multilingual improvements.
Our MaLA corpus’s classification system utilizes ISO codes to offer a standardized way to reference languages, reducing ambiguity, especially for languages with multiple names or dialects, and includes writing systems that further enhance precision, particularly for languages that use multiple scripts. This approach facilitates better data organization, model training, and cross-referencing with external linguistic datasets. Besides, we report per-language performance for some benchmarks in the Appendix, allowing readers to categorize the performance into different groups of languages as they wish. To avoid making the Appendix unnecessarily long, we will also share all results, including models’ generation, via GitHub, enabling the community to conduct further analyses with their preferred language taxonomy.
Q: Why do we not add which languages they officially support and separate the aggregated metrics for officially supported and unseen languages?
A: We can separate metrics for models that explicitly mention supported languages, but this is not possible for models like LLaMA and GEMMA 2, which do not provide such details. We report per-language performance for some benchmarks in the Appendix, allowing readers to categorize the performance into different groups of languages as they wish. To avoid making the Appendix unnecessarily long, we will also share all results, including models’ generation, via GitHub, enabling the community to conduct further analyses with their preferred metrics.
Appendix B Data Sources
B.1 Monolingual Data
B.1.1 List of Data Sources
The MaLA corpus harvests a wide range of datasets in multiple domains. Table 16 in the appendix lists the corpora and collections we used as monolingual data sources. The major ones include AfriBERTa (96), Bloom library (67), CC100 (27; 131) CulturaX (95), CulturaY (123), the Curse of Multilinguality (22), Evenki Life (71), Glot500 (50), GlotSparse (54), monoHPLT of HPLT v1.2 (31) from the HPLT project (14), Indigenous Languages Corpora (35), Indo4B (133), Lacuna Project (82), Languages of Russia (29), MADLAD-400 (61), Makerere Radio Speech Corpus (90), masakhane-ner1.0 (49), MC2 (139), mC4 (104), multilingual-data-peru (46), OSCAR 2301 (97) from the OSCAR project  [^24] https://oscar-project.org/, Tatoeba challenge monolingual collection (125), The Leipzig Corpora (42), Tigrinya Language Modeling (39), Wikipedia 20231101 (132) and Wikisource 20231201 (132). We exclude high-resource languages in CulturaX, HPLT, MADLAD-400, CC100, mC4, and OSCAR 2301. We exclude Gahuza and Pidgin in the AfriBERTa dataset. We filter out texts that mainly contain a date or timestamp in the Languages of Russia dataset. For Glot500-c, we filter out texts, which may come from train or test tests from datasets for machine translation, such as Flores200, Tatoeba, and mtdata. Despite the translation data being split into source and target languages in the Glot500-c corpus, we decide to filter them to avoid potential data leakage, especially since we use Flores200 as the evaluation benchmark.
@@ B.1.1 List of Data Sources
Glot500-c uses the following datasets: AI4Bharat, [^25] https://ai4bharat.org/ AIFORTHAI-LotusCorpus, [^26] https://github.com/korakot/corpus/releases/download/v1.0/AIFORTHAI-LotusCorpus.zip Add (37), AfriBERTa (96), AfroMAFT (5; 135), Anuvaad, [^27] https://github.com/project-anuvaad/anuvaad-parallel-corpus AraBench (109), AUTSHUMATO, [^28] https://autshumato.sourceforge.net/ Bloom (67), CC100 (27), CCNet (131), CMU_Haitian_Creole, [^29] http://www.speech.cs.cmu.edu/haitian/text/ CORP.NCHLT, [^30] https://repo.sadilar.org/handle/20.500.12185/7 Clarin, [^31] https://www.clarin.si/ DART (9), Earthlings (34), FFR, [^32] https://github.com/bonaventuredossou/ffr-v1/tree/master/FFR-Dataset Flores200 (30), GiossaMedia (44; 43), Glosses (20), Habibi (36), HinDialect (15), HornMT, [^33] https://github.com/asmelashteka/HornMT IITB (62), IndicNLP (93), Indiccorp (53), isiZulu, [^34] https://zenodo.org/record/5035171 JParaCrawl (86), KinyaSMT, [^35] https://github.com/pniyongabo/kinyarwandaSMT LeipzigData (42), Lindat, [^36] https://lindat.cz/faq-repository Lingala_Song_Lyrics, [^37] https://github.com/espoirMur/songs_lyrics_webscrap Lyrics, [^38] https://lyricstranslate.com/ MC4 (104), MTData (45), MaCoCu (18), Makerere MT Corpus, [^39] https://zenodo.org/record/5089560 Masakhane community, [^40] https://github.com/masakhane-io/masakhane-community Mburisano_Covid, [^41] https://repo.sadilar.org/handle/20.500.12185/536 Menyo20K (6), Minangkabau corpora (60), MoT (98), NLLB_seed (30), Nart/abkhaz, [^42] https://huggingface.co/datasets/Nart/abkhaz_text OPUS (124), OSCAR (118), ParaCrawl (17), Parallel Corpora for Ethiopian Languages (2), Phontron (94), QADI (3), Quechua-IIC (138), SLI_GalWeb.1.0 (8), Shami (4), Stanford NLP, [^43] https://nlp.stanford.edu/ StatMT, [^44] https://statmt.org/ TICO (11), TIL (84), Tatoeba, [^45] https://tatoeba.org/en/ TeDDi (85), Tilde (108), W2C (81), WAT (92), WikiMatrix (111), Wikipedia, [^46] https://huggingface.co/datasets/wikipedia Workshop on NER for South and South East Asian Languages (115), XLSum (47). We filter out Flores200 when processing the Glot500-c dataset. Glot500-c includes texts in languages, i.e., Azerbaijani, Gujarati, Igbo, Oromo, Rundi, Tigrinya and Yoruba, from the XLSum dataset.
B.2 Code
@@
Appendix C Additional Statistics of MaLA Corpus
C.1 Supported Languages
@@ C.2 Data Analysis
Appendix D Evaluation Setup
D.1 Prompts
D.1.1 Machine Translation
We use the prompt below for machine translation.
Translate the following sentence from {src_lang} to {tgt_lang}
[{src_lang}]: {src_sent}
[{tgt_lang}]:
D.1.2 Text Classification
The prompt template for SIB-200 is as follows:
{examples}
The topic of the news "\${text}" is
For Taxi-1500, the prompt template is as follows:
{examples}
The topic of the verse "\${text}" is
Appendix E Detailed Results
@@
Table 18: 0-shot results (ACC %) on XCOPA in all languages